# Towards gravimetry enhancement with squeezed states

### v1

@@ 0 $e^{-2r}$ $e^{2r}$
Conclusion. In this work, we have employed squeezed probe states to estimate the gravitational acceleration parameter. We analyze how both the amplitude and phase of the squeezing influence the sensitivity of the estimation. Our results show that the metrological advantage in gravitational acceleration estimation, compared to a vacuum probe state, is not solely determined by the squeezing strength but crucially depends on the squeezing phase, i.e, it depends on the direction of the applied squeezing in the probe state. Specifically, while probes squeezed along the canonical quadratures may fail to surpass the shot-noise limit, position–momentum correlated states with the same squeezing magnitude can exhibit a quantum Fisher information exceeding that of the vacuum state. Moreover, we identified optimal measurement strategies and squeezing phases that allow the classical Fisher information to reach the quantum limit. These findings deepen our understanding of how the phase of squeezed probes can be exploited to enhance precision in quantum gravimetry protocols.
@@
Acknowledgments. O.R.A. acknowledges the Graduate Program in Physics at Federal University of Piauí. L.S.M. acknowledges support from the National Institute of Science and Technology on the National Institute of Photonics (INFO) CNPq - INCT grant 409174/2024-6. J. F. G. S. acknowledges CNPq Grant No. 420549/2023-4, Fundect Grant No. 83/026.973/2024, and Universidade Federal da Grande Dourados for support. C.H.S.V. acknowledges the São Paulo Research Foundation (FAPESP) Grant No. 2023/13362-0 and Grant No. 2025/14546-2 for financial support and the Southern University of Science and Technology (SUSTech) for providing the workspace.

### v2

@@ 0 $e^{-2r}$ $e^{2r}$
Conclusion. In this work, we have employed squeezed probe states to estimate the gravitational acceleration parameter. We analyze how both the amplitude and phase of the squeezing affect the estimation sensitivity. Our results show that the metrological advantage in gravitational acceleration estimation, compared to a vacuum probe state, is not solely determined by the squeezing strength but crucially depends on the squeezing phase, i.e., the direction of the applied squeezing in the probe state. In particular, while probes squeezed along the canonical quadratures may fail to surpass the shot-noise limit, position–momentum correlated states with the same squeezing magnitude can exhibit a quantum Fisher information exceeding that of the vacuum state. Moreover, we identified optimal measurement strategies and squeezing phases that allow the classical Fisher information to reach the quantum limit. These findings deepen our understanding of how the phase of squeezed probes can be exploited to enhance precision in quantum gravimetry protocols. In realistic implementations, environmental noise and decoherence degrade both the squeezing strength and phase-space correlations, gradually reducing the QFI and eventually suppressing the quantum advantage. Although our analysis focuses on the ideal unitary regime, our results provide a fundamental benchmark for quantum-enhanced gravimetry and suggest that optimizing the squeezing phase may help mitigate decoherence-induced losses in practical scenarios.
@@
Acknowledgments. O.R.A. acknowledges the Graduate Program in Physics at Federal University of Piauí. L.S.M. acknowledges support from the National Institute of Science and Technology on the National Institute of Photonics (INFO) CNPq - INCT grant 409174/2024-6. J. F. G. S. acknowledges CNPq Grant No. 420549/2023-4, Fundect Grant No. 83/026.973/2024, and Universidade Federal da Grande Dourados for support. C.H.S.V. acknowledges the São Paulo Research Foundation (FAPESP) Grant No. 2023/13362-0 and Grant No. 2025/14546-2 for financial support and the Southern University of Science and Technology (SUSTech) for providing the workspace during the internship.